OpenStudy (anonymous):

Solve log(2x + 1) = 3

OpenStudy (mathstudent55):

Transform your equation with a log into the equation with an exponent. Use the definition of log below. \( \log_b x = y \iff b^y = x\)

OpenStudy (raden):

first, set 3 = log 1000 log (2x + 1) = 3 log (2x + 1) = log 1000 now take an equation : 2x + 1 = 1000 solve for x

OpenStudy (anonymous):

\[\log{(2x+1)}=3\] Use: \[\log_b x = y \iff b^y = x\] \[10^3 = (2x + 1)\] \[1000 = 2x +1\] \[999 = 2x\] \[999\div 2=199.5\] \[x=199.5\]